OpenStudy (anonymous):

@acxbox22 Find the solution to the following equation by transforming it into a perfect square trinomial. x2 – 12x = 13 A. {46, 55} B. {–1, 13} C. {–15, 27} D. {–21, 33}

OpenStudy (acxbox22):

you dont have to transform it into a perfect square trinomial subtract 13 on both sides x^2-12x-13=0 factor (x-13)(x+1)=0 x-13=0 x+1=0 x=13 x=-1 answer is B